OpenStudy (anonymous):

What is the first step in solving the equation 4.2y-5.2=18.5? Add 5.2 to each side of the equation. Subtract 5.2 from each side of the equation. Multiply each side of the equation by 4.2. Divide each side of the equation by 5.2.

OpenStudy (neer2890):

add 5.3 from each side of equation 4.2y-5.2+5.2=18.5+5.2 4.2y=23.7

OpenStudy (anonymous):

Cooper is measuring ingredients for his special dessert. He needs a total of 70.4 mL of milk and melted butter. The recipe calls for 30.6 mL of butter. How much milk, m, does Cooper need for his recipe? 2.3 mL 39.8 mL 40.2 mL 101 mL

OpenStudy (neer2890):

milk+butter= 70.4ml milk+30.6ml=70.4ml milk=(70.4-30.6)ml

OpenStudy (anonymous):

Sorry, what about this one? After getting 5 new rocks, George gave half of his rock collection to Susan. If George gave Susan 36 rocks, which equation could be used to determine how many rocks George started with?